### v2

# GRDD+: An Extended Greek Dialectal Dataset with Cross-Architecture Fine-tuning Evaluation

Abstract
We present an extended Greek Dialectal Dataset (GRDD+)  [^1] The full code for fine-tuning and the dataset GRDD+ are available at the following anonymous link: https://drive.google.com/drive/folders/1Xwfz08S8-9ZqMGd6EaNSje33LIaSE2E5?copy.that complements the existing GRDD dataset with more data from Cretan, Cypriot, Pontic and Northern Greek, while we add six new varieties: Greco-Corsican, Griko (Southern Italian Greek), Maniot, Heptanesian, Tsakonian, and Katharevusa Greek. The result is a dataset with total size 6,374,939 words and 10 varieties. This is the first dataset with such variation and size to date. We conduct a number of fine-tuning experiments to see the effect of good quality dialectal data on a number of LLMs. We fine-tune three model architectures (Llama-3-8B, Llama-3.1-8B, Krikri-8B) and compare the results to frontier models (Claude-3.7-Sonnet, Gemini-2.5, ChatGPT-5).
1.  Introduction
Modern Greek exhibits rich dialectal variation across different geographical regions. Despite this diversity, computational resources for these dialects remain limited, constraining the study and processing of regional linguistic varieties. Meanwhile, in the current rapid advancement of Natural Language Processing (NLP), Large Language Models (LLMs) have emerged at the forefront of research and development. However, LLMs frequently struggle with dialectal variations in lower-resourced languages, e.g. in Parts of Speech (POS) tagging and dialect identification faisal2025testing. While their dialect performance can surpass zero-shot transfer, it still falls behind the fine-tuned results faisal2024dialectbench. These limitations significantly impact their ability to generate contextually appropriate responses across regional dialects.
This paper introduces an extended dataset GRDD+ and fine-tuning experiments across multiple model architectures. Our study aims to evaluate how model adaptation can improve dialectal performance in Greek and provide new benchmarks for dialectal NLP. The remainder of the paper is organized as follows. Section 2 reviews related work. Section 3 describes the dataset and methodology, while Section GRDD+: An Extended Greek Dialectal Dataset with Cross-Architecture Fine-tuning Evaluation presents the fine-tuning experiments. Section GRDD+: An Extended Greek Dialectal Dataset with Cross-Architecture Fine-tuning Evaluation reports and discusses the results. Section GRDD+: An Extended Greek Dialectal Dataset with Cross-Architecture Fine-tuning Evaluation outlines future work, while Sections GRDD+: An Extended Greek Dialectal Dataset with Cross-Architecture Fine-tuning Evaluation and GRDD+: An Extended Greek Dialectal Dataset with Cross-Architecture Fine-tuning Evaluation discuss the limitations and present the conclusion and closing remarks, respectively.
2.  Related Work
Against this backdrop, resources for Modern Greek dialects remain scarce. Existing datasets include, among others, a small corpus for Griko anastasopoulos2018part, the Cypriot Greek version of the Multi-CAST corpus of annotated spoken texts hadjidas2015multi, and a database comprising 505 hours of recorded dialectal speech with linguistic and meta-linguistic annotations karasimos2008greed. To our knowledge, the GRDD corpus chatzikyriakidis2023grdd constitutes a first comprehensive effort to develop large-scale publicly available resources for Modern Greek dialects.
In parallel, several studies have emerged in the field of Greek computational dialectology, such as the development of Treebanks and parsers for Eastern Cretan in the framework of Universal Dependencies vakirtzian2025dialectal, the detection of Italian and Turkish loanwords in Greek dialects scherrerpsaltakichatzikyriakidis2025laonwords and computational analyses of the linguistic varieties of Cappadocian, Pharasiot, and Silliot bompolas2023computational. However, to the best of our knowledge, none of these efforts have attempted fine-tuning Large Language Models (LLMs) on Greek Dialectal Data.
3.  GRDD+ Dataset
3.1.  Collection Methodology
We focused on freely available dialectal data collected from the web. These include texts from blogs, websites, and publicly accessible literary sources such as songs, poems, folktales, dialogues and translations of works into the dialect by native speakers. Additionally, we collected dialectal data for certain varieties (Greco-Corsican, Griko, Heptanesian, Maniot and Pontic) from publicly available books using Optical Character Recognition (OCR) via Google Cloud Vision OCR  [^2] https://cloud.google.com/vision, subsequently removing all book metadata and retaining only the clean dialectal text. After data collection, we performed basic preprocessing on the data, including the removal of numbers, URLs, special characters, duplicate lines and extra white spaces.
Building upon the GRDD dataset chatzikyriakidis2023grdd, which consists of four dialects of Modern Greek, specifically Cretan, Pontic, Northern Greek, and Cypriot Greek, the present work seeks to extend and enhance the resource. Specifically, we enrich the existing dialectal corpora and incorporate six additional Greek dialectal varieties, as detailed below.
3.1.1.  Greco-Corsican
In the 1670s, Greek migrants from Mani settled in Cargèse, Corsica, forming a Greek-speaking community nicholas2005history. From the 1670s to the 1960s, a span of nearly three centuries, Greek was spoken in Cargèse, in relative isolation from other Greek-speaking communities. The variety, known as Greco-Corsican, has been the subject of detailed linguistic study phardys1888yleiskarifima; blanken1951grecs; parlangeli1952dom; rexine1966vayacacos. However, linguistic assimilation progressed rapidly, and by the 1930s only about 20 speakers of Greek remained. The language ultimately became extinct with the death of its last native speaker, Justine Voglimacci, in 1976.
3.1.2.  Griko (Southern Italian Greek)
Griko is a Greek dialect spoken in Grecìa Salentina, southern Italy, and recognized as a minority language. Officially, Grecìa Salentina consists of 12 villages: Calimera, Carpignano Salentino, Castrignano dei Greci, Corigliano d’Otranto, Cutrofiano, Martano, Martignano, Melpignano, Sogliano Cavour, Soleto, Sternatia and Zollino. Griko together with Grecanico of Calabria, form the endangered Italiot Greek group salminen1999unesco. Written in the Latin alphabet and only partly intelligible with Modern Greek, Griko now has fewer than 20,000 mostly elderly speakers chatzikyriakidis2010clitics.
3.1.3.  Heptanesian
Heptanesian is a Modern Greek dialect spoken on the Ionian Islands, including Corfu, Cephalonia, Lefkada, Zante, Ithaca, Kithira, Paxi and smaller islands such as Othoni, Antipaxi, and Antikithira kontosopoulos2000nikolaou. These islands were under Venetian rule from the late 14th to the late 18th century. Heptanesian exhibits Venetian and Italian influences primarily in its vocabulary, phonology (e.g.intonation), and morphology (e.g., the noun suffix –aδa < Ven –ADA), with syntax largely unaffected ralli2012verbal. Today, Heptanesian is gradually being abandoned in favor of Standard Modern Greek (SMG).
3.1.4.  Tsakonian
Tsakonian, a highly divergent modern form of Greek, still spoken in the eastern Peloponnese, is often considered distinct enough to be classified as a separate language from the rest of Modern Greek. As the only Modern Greek dialect that is not descended from the Hellenistic Koine, Tsakonian represents the main exception among Modern Greek varieties, deriving more or less directly from the ancient Doric dialect joseph1987modern; mackridge2010modern. horrocks2014greek refers to Tsakonian as a case of extreme dialectal resilience, exempt from the fundamental sound changes that shaped Modern Greek, such as the reversal of /u/ > /i/, while the dialect also exhibits numerous features that are unusual or unique compared to other Modern Greek varieties liosis2016tsakonian.
3.1.5.  Maniot
Maniot refers to the dialect spoken in the region of Laconian Mani. According to the traditional classification proposed by hatzidakis1892einleitung, which divides Modern Greek dialects into northern and southern groups, Maniot is categorized among the southern varieties. kontosopoulos2008dialects notes that Maniot constitutes a dialect distinct from the rest of the Peloponnesian varieties. The same view appears to be supported by trudgill2003modern, which emphasize the distinctiveness of the Maniot dialect.
The linguistic systems that appear to share similarities with Maniot include SMG, Cretan trudgill2003modern, Megarian, and, of course, several other Peloponnesian dialects pantelidis2001peloponnesiakos. pantelidis2001peloponnesiakos has argued that the similarities observed between SMG and the Peloponnesian dialects result from the influence of SMG on these dialects, and not vice versa, as had previously been claimed by mackridge1994neogreek; browning1969medieval; kontosopoulos2008dialects, among others.
3.1.6.  Katharevusa Greek
Katharevusa, described as a ‘ purist ’ (literally the purifying language) variety of SMG, served as the official written language of Greece from the establishment of the modern Greek nation-state until 1976 joseph1987modern; mackridge2010modern. This language variety was the middle solution during the language controversy  [^3] The language controversy, which originated in the 1760s, re-emerged with the establishment of the modern Greek state (1830) through the debate over which variety should serve as the official language of the newly independent nation mackridge2010modern. Katharevousa emerged as a kind of compromise between adopting Ancient Greek and the spoken form of SMG as the national language horrocks2014greek., and it was mostly used in written texts mackridge2010modern (for a contrasting view that argues that conditions of diglossia were developed between Katharevousa and SMG here: joseph1987modern). Katharevusa combined elements of both Ancient and Modern Greek, retaining much of the classical vocabulary and morphology while introducing intermediate forms such as ειµεϑα ` ωε αρε ´ ” ανδ ητον ` ηε/ςηε/ιτ ωας ´. Σψνταςτιςαλλψ, ιτ ωας ςλοςερ το ΣΜΓ, υςινγ ςονςτρυςτιονς λιϰε να + φινιτε ερβ ανδ τηε νεγατιε δεν, ψετ ιτ πρεςερεδ μανψ ανςιεντ παρτιςιπιαλ ςτρυςτυρες αβςεντ φρομ τηε ςποϰεν λανγυαγεmackridge2010modern· horrocks2014greek.
[…]
Δυρινγ psaltaki2025deixis’ς μαςτερ’ς τηεςις, ςηε ςτυδιεδ ῝ρεταν αδερβς εξπρεςςινγ δειξις. Ατ τηε τιμε, νο διαλεςταλ ςορπυς ωας ααιλαβλε, ςο ςηε ςρεατεδ α ςορπυς ςονταινινγ εξαμπλες οφ αδερβς δενοτινγ ηερε ανδ τηερε. Τηε ςορπυς ςομβινες τεξτς φρομ τηε ῝ρεταν Ρεναιςςανςε (15τη–17τη ς.) ςολλεςτεδ βψ Καϰλαμανις (2020) ανδ 62 φολϰλορε βοοϰς (1876–2020). Τηε ρεςυλτινγ ρεςουρςε, ῝ρετΔειΑδις psaltaki2025deixis, ις ιδεαλ φορ ρεςεαρςηερς ιντερεςτεδ ιν ῝ρεταν αδερβς, παρτιςυλαρλψ δειςτις εξπρεςςιονς, ανδ ις βεινγ οφφερεδ το τηε ρεςεαρςη ςομμυνιτψ φορ φυρτηερ ςτυδψ.
[…]
Τηε ΓΡΔΔ οριγιναλ ςορπυς ςομπριςες φουρ μαιν Γρεεϰ διαλεςτς: Ποντις Γρεεϰ, ῝ρεταν Γρεεϰ, ῝ψπριοτ Γρεεϰ, ανδ Νορτηερν Γρεεϰ. Ωε υςεδ τηε τερμ Ποντις Γρεεϰ το ρεφερ το τηε διαλεςτ ας ςποϰεν τοδαψ ιν μοδερν Γρεεςε, αλτηουγη α φορμ οφ Ποντις, Ρομεψϰα Ποντις, ις ςτιλλ ςποϰεν ιν ςομε ιλλαγες οφ Τραβζον ανδ ςυρρουνδινγ αρεας ιν πρεςεντ-δαψ Τυρϰεψ sitaridou2012cultural. ῝ρεταν Γρεεϰ ις ςποϰεν ον τηε ιςλανδ οφ ῝ρετε ανδ ις δεριεδ φρομ Κοινε Γρεεϰ mackridge1985moderngreek. ῝ψπριοτ Γρεεϰ ις ςποϰεν πριμαριλψ βψ Γρεεϰ ῝ψπριοτς, ας ωελλ ας βψ ςομε Τυρϰιςη ῝ψπριοτς. Τηε πρειους ερςιον οφ τηε ςορπυς ον Νορτηερν διαλεςτς ινςλυδεδ δατα ονλψ φρομ Κοζανι ανδ Γρεενα, βυτ ωε ηαε νοω εξτενδεδ ιτ το αλςο ινςλυδε Λεςβος, Σαμοτηραϰε ανδ Τηραςε ρεφλεςτινγ τηε βροαδερ ςςοπε οφ Νορτηερν διαλεςτς. Τηις ις ςομετηινγ ωορτη μεντιονινγ εεν τηουγη ωε ωιλλ νοτ διςςυςς τηε οριγιναλ ςορπυς διαλεςτς chatzikyriakidis2023grdd ιν δεταιλ.
Ωιτη τηε διαλεςτς οφ τηε ΓΡΔΔ οριγιναλ ςορπυς τηατ αρε ςηοων ιν ταβλε 1, τηε αδδιτιον οφ νεω δατα ρεςυλτς ιν ςυβςταντιαλ γροωτη αςροςς ςεεραλ διαλεςτς. Ποντις Γρεεϰ ινςρεαςες μοδερατελψ, ωιτη τηε νεω ωορδς ςοντριβυτινγ ρουγηλψ +8.1% το τηε οριγιναλ 867,935, φορ α τοταλ οφ 938,220 ωορδς. ῝ρεταν Γρεεϰ εξπεριενςες α μορε προνουνςεδ εξπανςιον, αδδινγ 583,808 ωορδς, αν ινςρεαςε οφ 64.8%, βρινγινγ ιτς τοταλ το 1,484,203 ωορδς. ῝ψπριοτ Γρεεϰ γροως μοδεςτλψ, ωιτη τηε νεω ωορδς ςοντριβυτινγ ρουγηλψ 2.1% το τηε οριγιναλ 1,345,849, φορ α τοταλ οφ 1,374,024 ωορδς. Νορτηερν Γρεεϰ, ινιτιαλλψ τηε ςμαλλεςτ οφ τηεςε διαλεςτς, μορε τηαν τριπλες ιν ςιζε ωιτη τηε νεω αδδιτιονς, ριςινγ βψ αβουτ 260.1% το ρεαςη 119,894 ωορδς. Τηις γροωτη ιμπροες δαταςετ ςοεραγε, ςυππορτινγ μορε ροβυςτ ςροςς-διαλεςταλ αναλψςες ανδ ςομπυτατιοναλ μοδελινγ.
Τηε νεωλψ αδδεδ δατα ινςλυδες ςεεραλ αριετιες τηατ ωερε νοτ πρεςεντ ιν τηε οριγιναλ δαταςετ. Κατηαρεουςα δομινατες ωιτη 1,515,982 ωορδς. Τςαϰονιαν αλςο ηας α ςυβςταντιαλ ρεπρεςεντατιον ωιτη 442,512 ωορδς, ηιγηλιγητινγ τηε εφφορτ το δοςυμεντ τηις ηιγηλψ ενδανγερεδ διαλεςτ. Γριςο, αν Ιταλο-Γρεεϰ μινοριτψ λανγυαγε, ςοντριβυτες 366,889 ωορδς, προιδινγ ιμπορταντ ςοεραγε φορ α Γρεεϰ αριετψ ουτςιδε Γρεεςε.
Σμαλλερ διαλεςτς ινςλυδε Ηεπτανεςιαν (50,311 ωορδς), Μανιοτ (30,692 ωορδς) ανδ Γρεςο-῝ορςιςαν (5,026 ωορδς). Τηερε ις αλςο ῝ρετΔειΑδ (47,186 ωορδς), ωηιςη ρεπρεςεντς α ςπεςιαλιζεδ ςυβςορπυς φοςυςινγ ον δειξις αδερβς ιν ῝ρεταν Γρεεϰ. Δεςπιτε τηειρ ςμαλλερ ςιζε, τηεςε αδδιτιονς αρε αλυαβλε φορ τηε πρεςερατιον ανδ αναλψςις οφ μινοριτψ ορ ρεγιοναλλψ ρεςτριςτεδ αριετιες ανδ φορ ςτυδιες οφ ςπεςιφις λινγυιςτις πηενομενα. Τηεςε αδδιτιονς ιμπροε τηε ςοεραγε οφ μινοριτψ, ρεγιοναλ ανδ ςπεςιαλιζεδ αριετιες, ςυππορτινγ μορε ροβυςτ ςροςς-διαλεςταλ αναλψζες ανδ ςομπυτατιοναλ μοδελινγ.Τηε οεραλλ ςιζε οφ τηε ςορπυς ηας αππροξιματελψ δουβλεδ φολλοωινγ τηε ινςορπορατιον οφ τηε νεω δατα. Ταβλε 1 ςηοως τηε διςτριβυτιον.
Διαλεςτ/῞αριετψ ΓΡΔΔ Ωορδ ςουντ Νεω Ωορδ ςουντ ΓΡΔΔ+ Ωορδ ῝ουντ
Ποντις 867,935 70,285 938,220
῝ρεταν 900,395 583,808 1,484,203
῝ψπριοτ 1,345,849 28,175 1,374,024
Νορτηερν 33,292 86,602 119,894
Κατηαρεουςα 1,515,982 1,515,982
Τςαϰονιαν 442,512 442,512
Γριςο 366,889 366,889
Ηεπτανεςιαν 50,311 50,311
Μανιοτ 30,692 30,692
Γρεςο-῝ορςιςαν 5,026 5,026
῝ρετΔειΑδ 47,186 47,186
Τοταλ 3,147,471 3,227,468 6,374,939
Ταβλε 1: Ωορδ ςουντς ιν τηε οριγιναλ ΓΡΔΔ ςορπυς, νεωλψ αδδεδ ωορδς φορ εξιςτινγ ανδ νεω διαλεςτς/αριετιες, ανδ τοταλ ωορδ ςουντς ιν ΓΡΔΔ+ περ διαλεςτ/αριετψ.
[…]
Ωε ςονςτρυςτεδ α διαλεςταλ φινε-τυνινγ δαταςετ φρομ ραω τεξτ ςορπορα ρεπρεςεντινγ φουρ Γρεεϰ ρεγιοναλ διαλεςτς φρομ τηε ΓΡΔΔ ςολλεςτιον: ῝ρεταν, Ποντις, Νορτηερν Γρεεϰ, ανδ ῝ψπριοτ Γρεεϰ. Το ςρεατε ςτρυςτυρεδ τραινινγ εξαμπλες φρομ τηε ραω τεξτ, ωε υςεδ α ςλιδινγ ωινδοω αππροαςη:
  1. 1.
    Τεξτ ις ςπλιτ ιντο ςηυνϰς οφ 100 ωορδς
  2. 2.
    ῝ηυνϰς ωιτη ατ λεαςτ 50 ωορδς αρε τυρνεδ ιντο προμπτ-ςομπλετιον παιρς:
    • Λονγερ ςηυνϰς ($\geq$ 80 ωορδς): Σπλιτ ιν ηαλφ, φιρςτ ηαλφ ις τηε προμπτ, ςεςονδ ηαλφ ις τηε ςομπλετιον
    • Σηορτερ ςηυνϰς (50-79 ωορδς): Τηε φυλλ ςηυνϰ ις τηε ςομπλετιον
  3. 3.
    Εαςη εξαμπλε ςταρτς ωιτη α διαλεςτ ινςτρυςτιον ιν Γρεεϰ (ε.γ., ᾽Γραψε ςτην ϰρητιϰη διαλεϰτο:᾽ φορ ῝ρεταν)
  4. 4.
    Ωε ρανδομλψ πιςϰ φρομ μυλτιπλε ινςτρυςτιον τεμπλατες περ διαλεςτ
Τηις γαε υς 26,118 τραινινγ εξαμπλες αςροςς αλλ φουρ διαλεςτς, ςαεδ ας ΘΣΟΝΛ φιλες. Ταβλε 2 ςηοως τηε διςτριβυτιον.
Φορ ῝ψπριοτ Γρεεϰ, ωε ςομβινεδ τωο ςεπαρατε ςορπορα: α ςυβςετ οφ ουρ πυβλιςλψ ααιλαβλε ςορπυς (5,625 εξαμπλες φρομ 562,522 ωορδς) ανδ τηε ΑΠΟαποιϰιοΠΟΙΗΣΗ ςορπυς (achilleos) (6,966 εξαμπλες φρομ 696,567 ωορδς), υςεδ ωιτη περμιςςιον φρομ τηε αυτηορς.
Διαλεςτ Ωορδς Εξαμπλες %
῝ρεταν 900,395 9,004 44.8%
Ποντις 418,997 4,190 20.8%
Νορτηερν 33,292 333 1.7%
῝ψπριοτ (πυβλις) 562,522 5,625 28.0%
῝ψπριοτ (ΑΠΟαποιϰιοΠΟΙΗΣΗ) 96,410 964 4.8%
Τοταλ 2,011,616 20,116 100%
Ταβλε 2: Φινε-τυνινγ δαταςετ διςτριβυτιον
Τηε διςτριβυτιον ρεφλεςτς τηε αρψινγ ααιλαβιλιτψ οφ ηιγη-χυαλιτψ διαλεςταλ ρεςουρςες ιν ΓΡΔΔ, ωιτη ῝ρεταν (44.8%) ανδ τηε ςομβινεδ ῝ψπριοτ δατα (32.8%) βεινγ ωελλ-ρεπρεςεντεδ, φολλοωεδ βψ Ποντις (20.8%) ανδ Νορτηερν Γρεεϰ (1.7%). Ωε πρεςερεδ τηις νατυραλ διςτριβυτιον το μαξιμιζε τηε υςε οφ ααιλαβλε διαλεςταλ δατα, τηουγη ωε αςϰνοωλεδγε τηις ιμβαλανςε ας α ποτεντιαλ λιμιτατιον τηατ μαψ αφφεςτ ρελατιε περφορμανςε αςροςς διαλεςτς.
[…]
Ωε φινε-τυνε τηρεε μοδελς:
  • Λλαμα-3-8Β: Μετα´ς ινςτρυςτιον-τυνεδ μυλτιλινγυαλ μοδελ
  • Λλαμα-3.1-8Β: Ενηανςεδ ερςιον ωιτη εξτενδεδ ςοντεξτ (128ϰ τοϰενς)
  • Κριϰρι-8Β: Γρεεϰ-ςπεςιαλιζεδ μοδελ βυιλτ ον Λλαμα-3.1-8Β, τραινεδ ον 56.7Β Γρεεϰ τοϰενς, τηε πρεμιερ ΛΛΜ φορ τηε Γρεεϰ λανγυαγε (krikri).
[…]
Ωε υςε ΛοΡΑ (hu2021lora) φορ εφφιςιεντ φινε-τυνινγ. Ταβλε 3 ςηοως ουρ ςεττινγς.
Παραμετερ ῞αλυε
ΛοΡΑ Ρανϰ ($r$) 16
ΛοΡΑ Αλπηα ($\alpha$) 32
ΛοΡΑ Δροπουτ 0.1
Ταργετ Μοδυλες χ_προϑ, ϰ_προϑ, _προϑ,
ο_προϑ, γατε_προϑ,
υπ_προϑ, δοων_προϑ
Τραιναβλε Παραμετερς $\sim$0.8% οφ βαςε μοδελ
Ταβλε 3: ΛοΡΑ ςονφιγυρατιον.
[…]
Ταβλε 4 ςηοως ουρ τραινινγ ηψπερπαραμετερς.
Ηψπερπαραμετερ ῞αλυε
Εποςης 3
Βατςη Σιζε περ Δειςε 2
Γραδιεντ Αςςυμυλατιον Στεπς 8
Εφφεςτιε Βατςη Σιζε 16
Λεαρνινγ Ρατε 3ε-4
ΛΡ Σςηεδυλερ ῝οςινε
Ωαρμυπ Στεπς 100
Οπτιμιζερ ΑδαμΩ
Ωειγητ Δεςαψ 0.01
Μαξ Γραδιεντ Νορμ 1.0
Πρεςιςιον βφλοατ16
Μαξ Σεχυενςε Λενγτη 512 τοϰενς
Ταβλε 4: Τραινινγ ηψπερπαραμετερς.
Αλλ εξπεριμεντς ραν ον ΑΩΣ μλ.π4δ.24ξλαργε ινςτανςες ωιτη Ν῞ΙΔΙΑ Α100 ΓΠϒς (40ΓΒ). Τραινινγ τοοϰ 4-6 ηουρς περ μοδελ, ωιτη πεαϰ μεμορψ υνδερ 35ΓΒ περ ΓΠϒ.
[…]
Ωε ςομπαρε ουρ τηρεε βαςε μοδελς, τηειρ τηρεε φινε-τυνεδ ερςιονς, ανδ τηρεε φροντιερ μοδελς, ῝λαυδε-3.7-Σοννετ, Γεμινι-2.5, ανδ ῝ηατΓΠΤ-5. Φορ εαςη διαλεςτ, ωε υςε 7 διφφερεντ προμπτς (ςηορτ ςτορψ, 3 μεδιυμ ςτοριες, λονγ ςτορψ, διαλογυε, ςρεατιε ωριτινγ), γιινγ 7 γενερατιονς περ μοδελ. Γιεν τηατ ωε ηαε α τοταλ οφ νινε μοδελς (3 φινε-τυνεδ + 3 βαςε μοδελς + 3 φροντιερ μοδελς), ωε ηαε 63 γενερατιονς περ διαλεςτ. Νατιε ςπεαϰερς εαλυατεδ τηε γενερατεδ τεξτς ον α 5-ποιντ ςςαλε ςηοων ιν (Ταβλε 5).
Σςορε Δεςςριπτιον
5 Απολυτα φυςιϰο - Νατιε-λεελ
4 Πολυ φυςιϰο - Μινορ ιςςυες
3 Μετρια φυςιϰο - Νοτιςεαβλε προβλεμς
2 Αφυςιϰο - Σιγνιφιςαντ προβλεμς
1 Εντελως αφυςιϰο - Νοτ διαλεςταλ
Ταβλε 5: Νατιε ςπεαϰερ εαλυατιον ςςαλε.
Μοδελ ῝ρεταν ῝ψπριοτ Ποντις Νορτηερν
Μεαν ΣΔ Μεαν ΣΔ Μεαν ΣΔ Μεαν ΣΔ
Λλαμα-3-8Β (βαςε) 1.15 0.49 1.52 1.03 1.11 0.52 1.32 0.89
Λλαμα-3-8Β (φινε-τυνεδ) 3.67 1.23 3.23 1.20 2.83 1.13 2.84 1.30
Λλαμα-3.1-8Β (βαςε) 1.13 0.45 1.38 0.78 1.00 0.00 1.30 0.92
Λλαμα-3.1-8Β (φινε-τυνεδ) 3.20 1.41 3.51 1.15 2.86 1.12 3.10 1.28
Κριϰρι-8Β (βαςε) 1.28 0.63 1.95 1.35 1.06 0.33 1.41 1.05
Κριϰρι-8Β (φινε-τυνεδ) 2.80 1.35 3.36 1.27 2.49 1.16 3.22 1.28
῝ηατΓΠΤ-5 2.49 1.37 3.36 1.08 2.14 0.96 3.54 0.89
῝λαυδε-3.7-Σοννετ 3.79 1.23 3.48 1.13 2.83 1.06 3.86 1.10
Γεμινι-2.5-Προ 1.63 0.92 2.47 1.20 1.06 0.33 2.02 1.00
Ταβλε 6: Νατιε ςπεαϰερ εαλυατιον ςςορες (1-5 ςςαλε) αςροςς διαλεςτς ανδ μοδελς. Μεαν ανδ ςτανδαρδ δειατιον (ΣΔ) ρεπορτεδ φορ εαςη διαλεςτ. ῝ρεταν ηαδ 16 ρατερς, ῝ψπριοτ 19 ρατερς, Νορτηερν 9 ρατερς ανδ Ποντις 5 ρατερς.
[…]
Τηε ρεςυλτς οφ ουρ εαλυατιον αρε ςηοων ιν Ταβλε 2. Ιντερ-ρατερ ρελιαβιλιτψ ωας αςςεςςεδ υςινγ μυλτιπλε μετριςς ανδ τηε ρεςυλτς αρε πρεςεντεδ ιν Ταβλε 7. Κριππενδορφφ´ς Αλπηα ρανγεδ φρομ 0.37 το 0.55 αςροςς διαλεςτς, ωηιςη ινδιςατες φαιρ το μοδερατε αγρεεμεντ ον αβςολυτε ςςορες. Ι῝῝(3,1), βαςιςαλλψ α τωο-ωαψ μιξεδ εφφεςτς μοδελ τρεατινγ ρατερς ας ρανδομ ανδ ιτεμς ας φιξεδ, ψιελδεδ αλυες βετωεεν 0.87 ανδ 0.96, δεμονςτρατινγ εξςελλεντ ςονςιςτενςψ ιν ρελατιε ρανϰινγς. Ωειγητεδ ῝οηεν´ς Καππα, ςαλςυλατεδ ας τηε αεραγε αςροςς αλλ ρατερ παιρς ανδ αςςουντινγ φορ ορδιναλ διςτανςε βετωεεν ρατινγς, ρανγεδ φρομ 0.39 το 0.54, φαλλινγ βετωεεν τηε οτηερ τωο μετριςς ιν ςενςιτιιτψ το αβςολυτε διφφερενςες.
Ιν τερμς οφ ινδιιδυαλ διαλεςτς, ῝ρεταν ςηοωεδ τηε ηιγηεςτ αγρεεμεντ αςροςς αλλ μετριςς (Κριππενδορφφ´ς $\alpha$=0.55, Ι῝῝(3,1)=0.96, ωειγητεδ $\kappa$=0.54). ῝ψπριοτ, ον τηε οτηερ ηανδ, ςηοωεδ τηε λοωεςτ Κριππενδορφφ´ς Αλπηα (0.37) ανδ ωειγητεδ Καππα (0.39), εεν τηουγη ιτ μαινταινς ηιγη ςονςιςτενςψ ιν ρελατιε ρανϰινγς (Ι῝῝(3,1)=0.95). Τηις παττερν μαψ ρεφλεςτ γρεατερ διαλεςταλ αριατιον ανδ/ορ αλςο ποιντ το τηε λαργερ, ανδ ποτεντιαλλψ μορε διερςε ρατερ ποολ φορ ῝ψπριοτ (19 ρατερς ερςυς 5-16 φορ οτηερ διαλεςτς). Ποντις ςηοωεδ ςυβςταντιαλλψ ηιγηερ εξαςτ αγρεεμεντ (39.7%), ωιτη ρατινγς αλςο ςλυςτερινγ ατ λοωερ ποιντς ιν τηε ςςαλε ςομπαρεδ το τηε οτηερ διελαςτς.
Τηε ηιγη Ι῝῝(3,1) αλυες δεμονςτρατε τηατ δεςπιτε διφφερενςες ιν ςςαλε υςαγε, ρατερς ςονςιςτεντλψ αγρεεδ ον ωηιςη τεξτς ωερε βεττερ ορ ωορςε, τηε ςριτιςαλ ρεχυιρεμεντ φορ αλιδ μοδελ ςομπαριςονς. Τηεςε ρεςυλτς αλιδατε τηε υςε οφ αεραγεδ ρατινγς ωηιλε αςϰνοωλεδγινγ τηε ινηερεντ ςυβϑεςτιιτψ ιν διαλεςταλ χυαλιτψ αςςεςςμεντ.
Μετρις Νορτηερν (9 ρατερς) ῝ρεταν (16 ρατερς) Ποντις (5 ρατερς) ῝ψπριοτ (19 ρατερς)
Κριππενδορφφ´ς $\alpha$ 0.429 0.545 0.425 0.373
Ι῝῝(2,1) 0.442 0.551 0.451 0.384
Ωειγητεδ $\kappa$ (αγ) 0.449 0.542 0.435 0.389
Εξαςτ αγρεεμεντ (%) 8.2 1.6 39.7 0.0
Ταβλε 7: Ιντερ-ρατερ ρελιαβιλιτψ αςροςς διαλεςτς. Κριππενδορφφ´ς $\alpha$, Ι῝῝(2,1), ανδ ωειγητεδ $\kappa$ αρε αππροπριατε φορ ορδιναλ ςςαλες ανδ ινδιςατε φαιρ το μοδερατε αγρεεμεντ (0.37–0.55). Εξαςτ αγρεεμεντ περςενταγες ςηοω εξπεςτεδ λοω αλυες φορ ςυβϑεςτιε μυλτι-ρατερ εαλυατιονς, εξςεπτ φορ Ποντις ωηιςη ηας α φαιρ εξαςτ αγρεεμεντ ςονςενςυς.
Τηερε αρε μανψ ιντερεςτινγ τηινγς το νοτε αβουτ τηε ρεςυλτς βοτη ιν τερμς οφ φινε-τυνινγ ας ωελλ ας μοδελ ςηοιςε. Τηε εαςιεςτ τηινγ το βε ςαιδ ις τηε ςομπαριςον οφ τηε βαςε ερςιονς οφ Λλαμα ανδ φροντιερ μοδελς λιϰε ΓΠΤ5, ῝λαυδε 3.7 ανδ Γεμινι 2.5-Προ. Λλαμα βαςε μοδελς ινςλυδινγ ϰριϰρι ηαε ςλοςε το ζερο διαλεςταλ ϰνοωλεδγε ωηιλε τηε φροντιερ μοδελς ρανγε ςεεμ το ποςςεςς διαλεςταλ ϰνοωλεδγε το αρψινγ δεγρεες, φρομ Γεμινι το ῝λαυδε.
Ανοτηερ διμενςιον ιν τηε διςςυςςιον ςονςερνς τηε ρελατιον βετωεεν Λλαμα βαςε ανδ φινε-τυνεδ ερςιονς. Ιτ ις ςλεαρ τηατ αλλ φινε-τυνεδ μοδελς αρε μυςη βεττερ τηαν τηε βαςε μοδελς, ςηοωινγ αν ινςρεαςε οφ βετωεεν 1.5-2 ποιντς αππροξιματελψ ιν τηειρ φινε-τυνεδ ερςιονς.
῝ομπαρινγ τηε φινε-τυνεδ μοδελς, ωε νοτιςε α νυμβερ οφ ιντερεςτινγ τηινγς βυτ νοτ αν αςροςς τηε βοαρδ ςλεαρ πιςτυρε. Τηε φιρςτ τηινγ τηατ ςτανδς ουτ ις τηατ Λλαμα-Κριϰρι ωηιςη ις τηε ονλψ μοδελ ωηιςη ις εξπλιςιτλψ τραινεδ ιν Μοδερν Γρεεϰ, δοες νοτ ςηοω τηε βεςτ περφορμανςε ουτ οφ τηε τηρεε. Λλαμα-ϰριϰρι ονλψ περφορμς βεττερ ιν τηε γενερατιον οφ Νορτηερν Γρεεϰ, ςςορες ςεςονδ φορ ῝ψπριοτ, ανδ τηιρδ (λαςτ) ιν τηε οτηερ τωο διαλεςτς, ι.ε. ῝ρεταν ανδ ῝ψπριοτ. Τηις μιγητ βε αν ινδιςατιον τηατ τηε οτηερ τωο μοδελς αρε μορε φλεξιβλε ιν λεαρνινγ τηε νεω αριετιες τηαν ϰριϰρι, δεςπιτε τηε λαττερ βεινγ εξπλιςιτλψ τραινεδ ον Μοδερν Γρεεϰ.
Φιναλλψ, ςομπαρινγ τηε φινε-τυνεδ 8Β ωιτη τηε τηρεε φροντιερ μοδελς, α νυμβερ οφ ιντερεςτινγ φινδινγς αλςο αριςε τηερε. Φιρςτ οφ αλλ, ῝λαυδε 3.7 ις ςονςιςτεντλψ ηιγη-περφορμινγ, τοππινγ τηε Νορτηερν ανδ ῝ρεταν ςατεγορψ, ανδ βεινγ ςεςονδ, ερψ ςλοςε το τηε φιρςτ, φορ ῝ψπριοτ ανδ Ποντις. Ιτ ις ιμπορταντ το νοτε ηερε τηατ τηε νεωερ ῝λαυδε ερςιονς (4 ονωαρδς) ηαε λοςτ τηειρ διαλεςταλ ςαπαβιλιτιες το ςομε εξτεντ, ανδ τηις ις ονε οφ τηε ρεαςονς τηατ ωε υςεδ τηις μοδελ ρατηερ τηαν τηε νεω ονες. Ωηατ τηις ηας ηαππενεδ ανδ το ωηατ εξτεντ, ις αν ιςςυε τηατ ωαρραντς μορε ινεςτιγατιον τηατ ωιλλ νοτ βε δονε ηερε. ΓΠΤ5 ις περφορμινγ δεςεντλψ ςονςιςτεντ, γιινγ χυιτε γοοδ περφορμανςες φορ ῝ψπριοτ ανδ Νορτηερν ανδ ρατηερ μεδιοςρε φορ τηε οτηερ τωο. Γεμινι ηας ςονςιςτεντλψ μεδιοςρε το ποορ περφορμανςε ρανγινγ φρομ 2.47 το 1.06.
Ιν τερμς οφ τηε ινδιιδυαλ διαλεςτς, ωε νοτιςε τηατ τηε ηιγηερ ςςορες αρε γιεν το Νορτηερν ανδ ῝ρεταν ρεςπεςτιελψ, φολλοωεδ βψ ῝ψπριοτ ανδ Ποντις. Αν ιντερεςτινγ χυεςτιον ηερε ςονςερνς ωηετηερ τηις ςλινε ηας ανψτηινγ το δο ωιτη τηε διςτανςε οφ τηεςε ινδιιδυαλ διαλεςτς το τηε δομιναντ αριετψ, Μοδερν Γρεεϰ, τηατ αλλ μοδελς ηαε ατ λεαςτ ςομε ϰνοωλεδγε οφ. Ιμπρεςςιονιςτις ιντυιτιονς αβουτ τηεςε διαλεςτς διςτατε τηατ ινδεεδ Νορτηερν ανδ ῝ρεταν αρε ςλοςερ το τηε δομιναντ αριετψ, ωηιλε ῝ψπριοτ ανδ λαςτλψ Ποντις αρε φαρτηερ αωαψ. [^4] In traditional Greek dialectology, there is a distinction between idioms and dialects. Basically, idioms were varieties that were closer to the dominant but not that far away to be considered dialects, and dialects varieties that were farther away from the dominant to be considered idioms. Northern and Cretan were usually considered idioms, Pontic and Cypriot dialects Οφ ςουρςε, τηε ιςςυε οφ λινγυιςτις διςτανςε ις λαργελψ υνεξπλορεδ ιν Γρεεϰ αριετιες, βυτ ιτ ωουλδ βε ιντερεςτινγ το ςεε ωηετηερ τηεςε ρεςυλτς ηερε, ςορρελατε ωιτη ςομε νοτιον οφ διςτανςε βετωεεν τηε δομιναντ αριετψ ανδ τηε ρεςπεςτιε διαλεςτς.
Λαςτλψ, τηε ρελατιονςηιπ βετωεεν τραινινγ δατα ςιζε ανδ μοδελ περφορμανςε αςροςς τηε φουρ διαλεςτς ις ςηοων το βε χυιτε ιντριγυινγ. ῝ρεταν ις τηε φιρςτ ιν ςιζε ωιτη 9,004 εξαμπλες (44.8%) ανδ περφορμς ωελλ (2.80-3.79), ωηιλε ῝ψπριοτ ηας 6,589 εξαμπλες (32.8% οφ τηε δαταςετ) ανδ αλλ τηρεε φινε-τυνεδ μοδελς ςςορινγ αβοε 3.0, ωιτη ονε φινε-τυνεδ μοδελ (Κριϰρι-8Β ατ 2.80) φαλλινγ βελοω τηε 3.0 τηρεςηολδ. Ποντις ηας 4,190 εξαμπλες (20.8%) βυτ ςονςιςτεντλψ ςςορες λοωεςτ (2.14-2.86), ωιτη αλλ τηρεε φινε-τυνεδ μοδελς φαιλινγ το ρεαςη 3.0. Συρπριςινγλψ, Νορτηερν Γρεεϰ ωιτη ονλψ 333 εξαμπλες (1.7%), μαναγες το αςηιεε ςτρονγ ςςορες (2.84-3.86), ωιτη ονλψ ονε φινε-τυνεδ μοδελ (Λλαμα-3-8Β ατ 2.84) βελοω 3.0. Τηε φαςτ τηατ ῝ψπριοτ ις τηε ονλψ διαλεςτ ωηερε αλλ φινε-τυνεδ μοδελς ςονςιςτεντλψ μαναγε το βρεαϰ τηε 3.0 μιγητ ινδιςατε τηε βενεφιτς οφ α λαργε δαταςετ τηατ ις τηε ςομβινατιον οφ τωο διερςε ςορπορα προιδινγ βεττερ ςοεραγε. Τηε Νορτηερν ρεςυλτς ρεμαιν νοταβλε, ας τηεψ δεςπιτε ηαινγ τηε λεαςτ δατα βψ φαρ, ιτ ματςηες ῝ρεταν´ς ςονςιςτενςψ βεττερ τηαν Ποντις δοες ωιτη 12 τιμες μορε τραινινγ εξαμπλες. Τηις παττερν μιγητ βε αν ινδιςατιον οφ λινγυιςτις διςτανςε φρομ Στανδαρδ Μοδερν Γρεεϰ, δατα χυαλιτψ διφφερενςες, ορ α ςομβινατιον οφ τηε τωο.
[…]
Ωιτη ρεςπεςτ το δαταςετ ςρεατιον, ωε ωουλδ λιϰε το δο α τηορουγη εαλυατιον οφ τηε δατα ςολλεςτεδ το ςομε υπ ωιτη ποτεντιαλλψ μορε φινε-γραινεδ ςατεγοριες. Φορ εξαμπλε, ῝ψπριοτ Γρεεϰ δατα ις ςομπριςεδ βψ δατα φρομ α νυμβερ οφ γενρες ινςλυδινγ βλογ ποςτς, λιτερατυρε ωριττεν ιν ῝ψπριοτ, τραδιτιοναλ ςονγς ανδ ριδδλες, ας ωελλ ας ςομε λιμιτεδ ςςιεντιφις τεξτς ιν ῝ψπριοτ. ῝λαςςιφψινγ ιντο μορε ςπεςιφις γενρες ωιλλ ποτεντιαλλψ ηελπ ρεςεαρςη ιν οτηερ φιελδς οφ Λινγυιςτιςς. Ον τηατ νοτε, ςοςιολινγυιςτις ςονςιδερατιονς ω.ρ.τ το τηε ςυρρεντ διγλοςςις ςιτυατιον ιν ῝ψπριοτ ας ωελλ ας τηε υςε οφ α ῝ψπριοτ Κοινε αρε αλςο ιςςυες τηατ μιγητ βε βενεφιτεδ φρομ ςομε οφ ουρ δατα, ιδεντιφψινγ παρτιςυλαρς παττερς τηατ αρε τηε ρεςυλτ οφ ςοδε-ςωιτςηινγ ορ οτηερ ςοςιαλινγυιςτιςαλλψ ρελεαντ μαρϰερς (ε.γ. Σηιβολλετη μαρϰερς οφ λοςαλ αριετιες (shibboleth)ξ).
Ονε οφ ουρ ιμμεδιατε πλανς οφ ςοντινυινγ τηις ωορϰ ςονςερνς τηε φινε-τυνινγ ον τηε ςιξ νεωλψ αδδεδ αριετιες (Γρεςο-῝ορςιςαν, Γριϰο, Ηεπτανεςιαν, Τςαϰονιαν, Μανιατιϰα, Κατηαρευςα). Τηις ωουλδ προιδε α μορε ςομπρεηενςιε ςοεραγε αςροςς αλλ ΓΡΔΔ+ διαλεςτς. Ωε αλςο πλαν το τεςτ α νυμβερ οφ αδδιτιοναλ αρςηιτεςτυρες (ε.γ. Μιςτραλ, Γεμμα) ανδ παραμετερ-εφφιςιεντ μετηοδς, ας ωελλ ας εξπλορινγ μυλτι-διαλεςτ μοδελς τηατ ηαε τηε αβιλιτψ το ηανδλε αλλ αριετιες ςιμυλτανεουςλψ.
Ωε πλαν το δεελοπ αυτοματις εαλυατιον μετριςς φορ διαλεςταλ χυαλιτψ ιν ορδερ το εναβλε ρεςεα. Ταςϰ-βαςεδ εαλυατιον (ςυμμαριζατιον, χυεςτιον-ανςωερινγ, τρανςλατιον) ςαν ςομπλεμεντ ουρ γενερατιον αππροαςη, ωηιλε, ατ τηε ςαμε τιμε, προιδε ινςιγητς ιντο διαλεςταλ ςομπρεηενςιον ςαπαβιλιτιες.
Λαςτλψ, ωε πλαν το ςοντινυε εξπανδινγ τηε ΓΡΔΔ δαταςετ βψ ενηανςινγ βοτη τηε ςυρρεντ διαλεςταλ ςορπορα ανδ αδδινγ νεω διαλεςτς το βροαδεν ιτς ςοεραγε.
[…]
Δαταςετ ιμβαλανςε. Τηε δαταςετ υςεδ φορ φινε-τυνινγ ις ερψ ιμβαλανςεδ. ῝ρεταν ςομπριςες 44.8% (9,004 εξαμπλες), ωηιλε Νορτηερν Γρεεϰ ονλψ 1.7% (333 εξαμπλες). Οφ ςουρςε, τηις ρεφλεςτς τηε αρψινγ ρεςουρςε ααιλαβιλιτψ, ανδ ιτ ις υνδερςτανδαβλε το ςομε εξτεντ, βυτ ςαν αφφεςτ τηε ρελατιε περφορμανςε αςροςς διαλεςτς.
Εαλυατιον ςυβϑεςτιιτψ. Τηε ρεςυλτς ωε ηαε ςηοω μοδερατε αγρεεμεντ λεελς (Κριππενδορφφ´ς $\alpha$=0.37–0.55). Τηις ρεφλεςτς αν ινηερεντ ςυβϑεςτιιτψ ιν τηεςε τψπε οφ νατυραλνεςς ϑυδγμεντς. Ι῝῝(3,1) αλυες (0.87–0.96) ινδιςατε εξςελλεντ ςονςιςτενςψ ιν ρελατιε ρανϰινγς, βυτ, ηοωεερ, τηε ρατινγ αριαβιλιτψ μιγητ ςυγγεςτ α νεεδ φορ μορε ςτρυςτυρεδ εαλυατιον προτοςολς ορ εεν ρατερ τραινινγ. Λαςτλψ, ουρ εαλυατιον υςες ονλψ 7 προμπτς περ διαλεςτ, πριμαριλψ ναρρατιε ταςϰς, ανδ ας ςυςη δο νοτ ςαπτυρε α φυλλερ ρανγε οφ υςαγε ςοντεξτς.
Λιμιτεδ ςςοπε. Ωε φινε-τυνεδ τηρεε 8Β μοδελς ωιτη α ςινγλε ΛοΡΑ ςονφιγυρατιον ανδ τηεν ςομπαρεδ τηεμ αγαινςτ τηρεε φροντιερ μοδελς. Φυρτηερμορε, τηε υνεξπεςτεδ υνδερπερφορμανςε οφ Κριϰρι-8Β, δεςπιτε ιτς Γρεεϰ-ςπεςιφις τραινινγ, μεριτς φυρτηερ ινεςτιγατιον.
Λινγυιςτις διςτανςε. Ουρ διςςυςςιον οφ περφορμανςε ρελατιε το λινγυιςτις διςτανςε φρομ Στανδαρδ Μοδερν Γρεεϰ ρεμαινς χυαλιτατιε. Γρεεϰ διαλεςτολογψ λαςϰς ςτανδαρδιζεδ διςτανςε μετριςς τηατ ωουλδ εναβλε ςψςτεματις τεςτινγ οφ τηις ρελατιονςηιπ.
Σοςιολινγυιςτις φαςτορς. Ωε δο νοτ αςςουντ φορ διγλοςςια, ςοδε-ςωιτςηινγ, ρεγιςτερ, γενρε, ορ ωιτηιν-διαλεςτ αριατιον (ε.γ., λοςαλ αριετιες ς. ῝ψπριοτ Κοινε (tsiplakou)).
[…]
Ωε πρεςεντεδ ΓΡΔΔ+, αν εξτενδεδ Γρεεϰ διαλεςταλ δαταςετ τηατ ινςλυδες δατα φρομ 10 αριετιες οφ Γρεεϰ. 4 οφ τηε διαλεςτς ωερε παρτ οφ τηε εξιςτινγ ΓΡΔΔ δαταςετ (῝ρεταν, ῝ψπριοτ, Ποντις, Νορτηερν) ανδ ηαε βεεν εξπανδεδ ιν τερμς οφ ςοεραγε, ανδ ςιξ νεω αριετιες (Γρεςο-῝ορςιςαν, Γριϰο, Ηεπτανεςιαν, Τςαϰονιαν, Μανιοτ, ανδ Κατηαρευςα) ωερε αδδεδ. Τηις ις α γοοδ βαςις φορ α ςομπρεηενςιε ρεςουρςε φορ Γρεεϰ διαλεςταλ ΝΛΠ.
Ωε τηεν εξπεριμεντεδ ωιτη α νυμβερ οφ φινε-τυνινγ εξπεριμεντς υςινγ τηρεε 8Β παραμετερ μοδελς (Λλαμα-3, Λλαμα-3.1, ανδ Κριϰρι). Τηε ρεςυλτς ςηοω τηατ ταργετεδ διαλεςταλ φινε-τυνινγ ιμπροες γενερατιον χυαλιτψ, ςηοωινγ γαινς οφ 1.5-2 ποιντς ον α 5 ποιντ ςςαλε οφ διαλεςταλ νατυραλνεςς. Τηις ις εεν τρυε ωιτη ρελατιελψ μοδεςτ αμουντς οφ τραινινγ δατα, ας ειδενςεδ βψ Νορτηερν Γρεεϰ, τηατ αςηιεςες ςτρονγ περφορμανςε (2.84-3.86) δεςπιτε ηαινγ ονλψ 333 τραινινγ εξαμπλες.
῝ομπαριςον ωιτη φροντιερ μοδελς ςηοως α ρατηερ νυανςεδ περφορμανςε πιςτυρε. ῝λαυδε-3.7-Σοννετ αςηιεες τηε ηιγηεςτ ςςορες φορ ῝ρεταν (3.79) ανδ Νορτηερν Γρεεϰ (3.86), ωηιλε τηε φινε-τυνεδ Λλαμα-3.1-8Β ουτπερφορμς αλλ μοδελς ον ῝ψπριοτ (3.51) ανδ Ποντις (2.86). Τηις ις αν ινδιςατιον τηατ ςπεςιαλιζεδ φινε-τυνινγ ςαν εναβλε ςμαλλερ μοδελς το εξςεεδ φροντιερ μοδελ περφορμανςε ον ςπεςιφις διαλεςτς. ῝ηατΓΠΤ-5 ςηοως ςολιδ, αλβειτ ινςονςιςτεντ περφορμανςε αςροςς διαλεςτς, ανδ Γεμινι-2.5-Προ ις υνδερπερφορμινγ αςροςς αλλ διαλεςτα. Νοταβλψ, βαςε Λλαμα μοδελς (ινςλυδινγ τηε Γρεεϰ-ςπεςιαλιζεδ Κριϰρι) ςηοω νεαρ-ζερο διαλεςταλ ςαπαβιλιτιες, ηιγηλιγητινγ τηε ςριτιςαλ ιμπορτανςε οφ διαλεςταλ τραινινγ δατα.
Α νυμβερ οφ οτηερ φινδινγς ωαρραντ φυρτηερ ινεςτιγατιον: (1) Κριϰρι-8Β φινε-τυνεδ ις υνδερπερφορμινγ ρελατιε το μυλτιλινγυαλ Λλαμα μοδελς δεςπιτε ιτς Γρεεϰ-ςπεςιφις τραινινγ, (2) τηερε ις α νον-λινεαρ ρελατιονςηιπ βετωεεν τραινινγ δατα ςιζε ανδ περφορμανςε (Νορτηερν ουτπερφορμινγ Ποντις δεςπιτε ηαινγ 12 τιμες λεςς δατα), ανδ (3) τηε ςορρελατιον βετωεεν διαλεςταλ περφορμανςε ανδ λινγυιςτις διςτανςε φρομ Στανδαρδ Μοδερν Γρεεϰ.
Ωε βελιεε τηατ τηις δαταςετ ανδ ουρ φινδινγς ςαν φυνςτιον ας α ςολιδ φουνδατιον φορ φυτυρε ωορϰ ον Γρεεϰ διαλεςταλ ΝΛΠ, ςινςε ωε ςηοω τηατ εεν ςμαλλ αμουντς οφ ηιγη-χυαλιτψ διαλεςταλ δατα ςαν εναβλε εφφεςτιε φινε-τυνινγ. Ωε ρεαλλψ ηοπε τηατ τηις ρεςουρςε ωιλλ εναβλε ρεςεαρςη νοτ ονλψ ιν ΝΛΠ βυτ αλςο ιν ςοςιολινγυιςτιςς, διαλεςτολογψ, ανδ λανγυαγε δοςυμεντατιον φορ Γρεεϰ ανδ οτηερ λανγυαγες ωιτη ριςη διαλεςταλ αριατιον.
[…]
Ωε τηανϰ Ανδρι Αςηιλλεος, Σπψρος Αρμοςτις, ανδ Ελενα Σοϰρατους φορ γραντινγ περμιςςιον το υςε δατα φρομ τηε ΑΠΟαποιϰιοΠΟΙΗΣΗ ςορπυς. Ωε αλςο τηανϰ Πανος Μαρνερις φορ γιινγ υς περμιςςιον το ςςραπε ανδ υςε τηε Τςαϰονιϰα δατα φουνδ ιν ηις ωεβςιτε. Εροφιλι Πςαλταϰι ρεςειεδ φυνδινγ φρομ τηε Ευροπεαν ϒνιον’ς Ηοριζον Ευροπε ρεςεαρςη ανδ ιννοατιον προγραμ υνδερ τηε Μαριε Σϰłοδοωςϰα-῝υριε γραντ αγρεεμεντ Νο 101177564—ΗΑΙΦ. ῝ο-φυνδεδ βψ τηε Ευροπεαν ϒνιον. ῞ιεως ανδ οπινιονς εξπρεςςεδ αρε ηοωεερ τηοςε οφ τηε αυτηορ(ς) ονλψ ανδ δο νοτ νεςεςςαριλψ ρεφλεςτ τηοςε οφ τηε Ευροπεαν ϒνιον ορ τηε Ευροπεαν Ρεςεαρςη Εξεςυτιε Αγενςψ (ΡΕΑ). Νειτηερ τηε Ευροπεαν ϒνιον νορ τηε γραντινγ αυτηοριτψ ςαν βε ηελδ ρεςπονςιβλε φορ τηεμ. Στεργιος ῝ηατζιϰψριαϰιδις γρατεφυλλψ αςϰνοωλεδγες φυνδινγ φρομ Αμαζον (προϑεςτ: Νευραλ-Σψμβολις Ιντεγρατιον φορ Ενηανςεδ Νατυραλ Λανγυαγε Προςεςςινγ (ΝΙΕΛΣ)) τηατ προιδεδ ςομπυτατιοναλ ςυππορτ φορ τηε φινε-τυνινγ εξπεριμεντς δεςςριβεδ ιν τηε παπερ. Στεργιος ῝ηατζιϰψριαϰιδις ις αλςο παρτιαλλψ φυνδεδ βψ τηε Ευροπεαν ϒνιον (ΕΡ῝ ΑΔΓ, ΠηψλΠροΓραμμ, 101096554). ῞ιεως ανδ οπινιονς εξπρεςςεδ αρε ηοωεερ τηοςε οφ τηε αυτηορ(ς) ονλψ ανδ δο νοτ νεςεςςαριλψ ρεφλεςτ τηοςε οφ τηε Ευροπεαν ϒνιον ορ τηε Ευροπεαν Ρεςεαρςη ῝ουνςιλ Εξεςυτιε Αγενςψ. Νειτηερ τηε Ευροπεαν ϒνιον νορ τηε γραντινγ αυτηοριτψ ςαν βε ηελδ ρεςπονςιβλε φορ τηεμ.